a father is now four times as old as his son. five years ago he was seven times as old as his son. how old are tehy now?

Answer by stanbon(75887)

a father is now four times as old as his son. five years ago he was seven times as old as his son. how old are they now?
-----------------------
Equations:
f = 4s
f-5 = 7(s-5)
----
Substitute for "f" and solve for "s":
4s - 5 = 7s - 35
3s = 30
s = 10 (Son's age now)
----
solve for "f";
f = 4s
f = 40 (Father's age now)
==============================
Cheers,
Stan H.
==============================